OpenStudy (anonymous):

x is a variable such that if 20% of it is added to its fifth the result is equal to 12 subtracted from seven tenths of x. Find x.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

2/5x = 7/10x - 12 4/10x = 7/10x - 120/10 4x = 7x - 120 3x = 120 x = 40

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you